OpenStudy (anonymous):

Show all your work to receive full credit for this problem. Solve the system of equations using the Addition method. 2a + 3b = -1 3a + 5b = -2

OpenStudy (anonymous):

2a + 3b = -1 multiply by 3 3a + 5b = -2 multiply by -2 we get 6a + 9b = -3 -6a - 10b = 4 add the above two, to get - b = 1 so b = -1 now substitute b = -1 in first given equation ie. 2a + 3b = -1 2a + 3(-1) = -1 2a - 3 = -1 2a = -1 + 3 a = 2/2 so a = 1 solutions set (a, b) = (1, -1)